OpenStudy (gabylovesyou):

simplify 12f^6g / 9f^3

OpenStudy (anonymous):

\[\frac{12f^{6g}}{9f^3}\]Is that the problem?

OpenStudy (gabylovesyou):

yes :p

OpenStudy (anonymous):

its like a^m/a^n = a^(m+n)

OpenStudy (gabylovesyou):

what would M and n be ?

OpenStudy (anonymous):

6g and 3

OpenStudy (anonymous):

12/9 is simply 4/3 so the answer you've got is now simplified. hope u understood

OpenStudy (gabylovesyou):

kinda... whats the answer all together

OpenStudy (anonymous):

4/3 ( f^(6g-3))

OpenStudy (gabylovesyou):